# Toward Attention-based TinyML: A Heterogeneous Accelerated Architecture and Automated Deployment Flow

### v1

@@ VI Conclusion
We presented a flexible hardware architecture template, enabling collaborative accelerated execution of emerging Attention-based workloads that can be easily extended for the demands of future networks. By integrating our hardware template in Deeploy, we demonstrate a flexible deployment flow capable of efficiently mapping both accelerator-specific and generic DNN operators on our target architecture. We demonstrate the first E2E deployment of an encoder transformer network, MobileBERT, on a parallel heterogeneous accelerator enhanced MCU. Our implementation, which leverages ITA for computing the MHA and Linear layers and the cluster cores for auxiliary operators, achieves state-of-the-art throughput of with an energy efficiency of at and .

### v2

@@ VI Conclusion
We presented a flexible hardware-software architecture template, enabling collaborative accelerated execution of emerging Attention-based workloads that can be easily extended for the demands of future networks. By integrating our hardware template in Deeploy, we demonstrate a flexible deployment flow capable of efficiently mapping both accelerator-specific and generic DNN operators on our target architecture. We demonstrate the first E2E deployment of multiple Transformer-based encoder models on a parallel heterogeneous accelerator-enhanced MCU. Our implementation, which leverages ITA for computing the MHA and Linear layers and the cluster cores for auxiliary operators, achieves state-of-the-art throughput of with an energy efficiency of . This enables inference rates of at for MobileBERT, at for DINOv2-Small, and at for encoder block of Whisper.